Please help me solve this equation
solve by elimination method.
0.3x-0.2y=4
0.5x+0.3y=71/23
The solution is _____ (type an ordered pair)

Answer by Alan3354(69443)

solve by elimination method.
0.3x-0.2y=4 Eqn 1
0.5x+0.3y=71/23 Eqn 2
------------
9x - 6y = 120 Eqn 1 times 30
10x+ 6y = 1420/23 Eqn 2 times 20
-----------------Add
19x = 4180/23
x = 220/23
------------
2200/23 + 6y = 1420/23
6y = -780/23
y = -130/23